# Sparse Model Soups: A Recipe for Improved Pruning via Model Averaging

### v1

@@ 5 Discussion
Efficient, high-performing sparse networks are crucial in resource-constrained environments. However, sparse models cannot easily leverage the benefits of parameter averaging. We addressed this issue proposing SMS, a technique that merges models while preserving sparsity, substantially enhancing IMP and outperforming multiple baselines. By integrating SMS into magnitude-pruning methods during training, we elevated their performance and competitiveness. Despite the focus on pruning, a single type of network compression, we think that our work serves as an important step towards understanding and improving sparsification algorithms. A potential limitation of SMS may be the computational cost, demanding multiple, yet parallelizable, training runs, which might be restrictive in some settings.
@@ Appendix A Technical details and training settings
Training settings and metrics used.
Table 3 shows the exact pretraining settings for each dataset-architecture pair, reporting the number of epochs used for pretraining, the batch size, weight decay as well as the learning rate used. We kept momentum at the default value of 0.9. The last column reports the performance we achieve when performing regular dense training. We stick to SGD as the optimizer, noting that a variety of other optimization methods for training deep neural networks exist [see e.g. Kingma and Ba, 2014, Pokutta et al., 2020]. For image classification tasks, we report the top-1 test accuracy being the fraction of correctly classified test samples. For semantic segmentation, we used pretrained backbones and evaluated the mean Intersection-over-Union (IoU) on the validation dataset. For NMT, we report the BLEU score on the test set with sequence length limited to 128. We utilized label smoothing and gradient clipping for MaxViT. If needed, we report the theoretical speedup, a metric indicating the FLOPs ratio for inference between dense and sparse models [Blalock et al., 2020]. The speedup, defined as $F_{d}/F_{s}$ where $F_{d}$ and $F_{s}$ are the FLOPs required for dense and pruned models respectively, depends solely on the distribution of pruned weights, not on the values attained by non-zero parameters. FLOPs are computed using a single test set batch, with code adapted from the ShrinkBench framework [Blalock et al., 2020].
Table 3: Exact training configurations in our experiments.
@@
Identifying which weights to remove is essential for successful magnitude pruning, with multiple methods developed to address this. Zhu and Gupta [2017] presented the Uniform allocation that prunes each layer to the same relative sparsity level. Gale et al. [2019] refined this approach to Uniform+, leaving the first convolutional layer dense and capping pruning in the final fully-connected layer at 80%. Evci et al. [2020] reformulate the Erdős-Rényi kernel (ERK) [Mocanu et al., 2018] to consider layer and kernel dimensions for layerwise sparsity distribution. Lee et al. [2020] suggested Layer-Adaptive Magnitude-based Pruning (LAMP), which minimizes output distortion at pruning time measured by the $L_{2}$-distortion on the worst-case input. Throughout this work, we stick to the Global allocation, in which all trainable parameters treated as a single vector and a global threshold is computed to remove parameters independent of the layer they belong to. In experiments where we prune convolutional filters instead of weights, we adopt the $L_{2}$-norm criterion from Li et al. [2016], ensuring a uniform sparsity distribution across layers.
@@
We follow the recommendations of Evci et al. [2020] and Dettmers and Zettlemoyer [2019] and refrain from pruning biases and batch-normalization parameters, as their negligible weight contribution is offset by their significant performance impact. Moreover, for GMP experiments, we opt for the global selection criterion due to its superior performance compared to Uniform+ [Zimmer et al., 2023].
@@
Throughout our experiments, we also explored variants of LearnedSoup [Wortsman et al., 2022a], which learns the coefficients $\lambda_{i}$ to maximize the validation accuracy. Specifically, we observed improvements when utilizing knowledge distillation techniques [Hinton et al., 2015] with the original pretrained model as the teacher, instead of minimizing the validation loss as suggested by Wortsman et al. [2022b]. Nevertheless, these improvements were marginal, so we opted for UniformSoup and GreedySoup for simplicity and to avoid introducing new hyperparameters.
@@
Appendix B Full results
This section contains additional tables and plots. The subsections follow the same structure as the main experimental section.
@@ B.1 Evaluating Sparse Model Soups
Table 4: WideResNet-20 on CIFAR-100 (unstructured pruning): Test accuracy comparison of SMS to several baselines for target sparsities 90% (top) and 98% (bottom) given three prune-retrain cycles. We report results using UniformSoup as well as GreedySoup for merging. Results are averaged over multiple seeds with standard deviation included. The best value is highlighted in bold.
@@
Table 5: ResNet-18 on CIFAR-10 (unstructured pruning): Test accuracy comparison of SMS to several baselines for target sparsity 98% given three prune-retrain cycles. We report results using UniformSoup as well as GreedySoup for merging. Results are averaged over multiple seeds with standard deviation included. The best value is highlighted in bold.
@@
Table 6: ResNet-18 on CIFAR-10 (unstructured pruning): Test accuracy comparison of SMS to several baselines for target sparsities 80%, 90%, 95% in the One Shot setting and a retrain length of 20 epochs. We report results using UniformSoup as well as GreedySoup for merging. Results are averaged over multiple seeds with standard deviation included. The best value is highlighted in bold.
@@
Table 7: MaxViT on ImageNet (unstructured pruning): Test accuracy comparison of SMS to several baselines for target sparsities 75%, 80%, 85% in the One Shot setting and a retrain length of 10 epochs. We report results using UniformSoup as well as GreedySoup for merging. Results are averaged over multiple seeds with standard deviation included. The best value is highlighted in bold.
@@
Table 8: PSPNet on Cityscapes (unstructured pruning): Validation accuracy comparison of SMS to several baselines for target sparsity 90% given two prune-retrain cycles. We only report results for UniformSoup since we evaluate on the validation dataset. Results are averaged over multiple seeds with standard deviation included. The best value is highlighted in bold.
@@
SMS (uniform) 58.61 ±0.22 58.96 ±0.35 59.28 ±0.73 59.01 ±0.35
best candidate 58.65 ±0.69 58.91 ±0.56 58.83 ±0.54 58.65 ±0.34
mean candidate 58.33 ±0.28 58.35 ±0.30 58.47 ±0.23 58.28 ±0.36
IMP+ 58.76 ±0.18 58.37 ±0.53 58.52 ±0.39 58.63 ±0.38
IMP-AVG — N/A — 58.79 ±0.40 59.01 ±0.07
IMP — 58.40 ±0.28 — — 58.41 ±0.09 —
Table 9: PSPNet on Cityscapes (unstructured pruning): Validation accuracy comparison of SMS to several baselines for target sparsities 60%, 70%, 80% and 90% in the One Shot setting and a retrain length of 10 epochs. We only report results for UniformSoup since we evaluate on the validation dataset. Results are averaged over multiple seeds with standard deviation included. The best value is highlighted in bold.
@@
SMS (uniform) 58.85 ±0.46 58.04 ±0.84 57.80 ±0.20 57.61 ±0.51
best candidate 58.61 ±0.86 58.00 ±0.84 57.37 ±0.14 57.45 ±0.44
mean candidate 58.17 ±0.69 57.69 ±0.72 57.21 ±0.16 56.91 ±0.35
IMP$+$ 58.08 ±0.03 58.41 ±0.22 57.69 ±0.23 57.36 ±0.06
IMP 58.09 ±0.57 57.59 ±0.62 57.42 ±0.44 56.94 ±0.54
Table 10: T5 on WMT16 (unstructured pruning): Test accuracy comparison of SMS to several baselines for target sparsities 50%, 60%, 70% in the One Shot setting. We report results using UniformSoup merging. Results are averaged over multiple seeds with standard deviation included. The best value is highlighted in bold.
@@
IMP+ 25.36 ±0.12 25.09 ±0.05 24.00 ±0.04
@@
Table 11: WideResNet-20 on CIFAR-100 (structured pruning): Test accuracy comparison of SMS to several baselines for target sparsities 60% (top) and 80% (bottom) given three prune-retrain cycles. We report results using UniformSoup as well as GreedySoup for merging. Results are averaged over multiple seeds with standard deviation included. The best value is highlighted in bold.
@@
Table 12: ResNet-18 on CIFAR-10 (structured pruning): Test accuracy comparison of SMS to several baselines for target sparsities 40%, 50%, 60% in the One Shot setting and a retrain length of 20 epochs. We report results using UniformSoup as well as GreedySoup for merging. Results are averaged over multiple seeds with standard deviation included. The best value is highlighted in bold.
@@ B.2 Examining Sparse Model Merging
Exploring parameters beyond random seeds.
@@
Exploring limitations: Instability to randomness and recovering it.
@@
Figure 6: ResNet-18 on CIFAR-10: (a) Accuracy difference between the soup and best performing model after One Shot pruning and retraining. The lines for UniformSoup and GreedySoup show the envelope considering all retraining schedules and durations. (b) Difference between the soup and prolonged IMP version after One Shot pruning to 90% and 98% sparsity. The x-axis shows the retraining length of the $m=3$ models in the soup, while the prolonged IMP version is retrained three times as long. Results are averaged over multiple random seeds with min-max bands indicated.
Reducing compression-induced unfairness.
@@
Classification model performance, usually measured by the top-1 accuracy, can mask the disproportionate effect of compression on individual class performance [Hooker et al., 2019, 2020, Paganini, 2020]. Pruning often sacrifices difficult samples, benefiting well-performing classes and worsening the performance of others [Tran et al., 2022]. Recent research by Ko et al. [2023] highlights the benefits of prediction ensembling in enhancing fairness metrics, including minority group performance.
@@
Table 13: ResNet-18 on Celeb-A: Comparison of the pretrained (i.e. dense) base model against SMS and IMP for different sparsity levels. We indicate the top-1 test accuracy as well as the recall on the four different sub-groups. All results are averaged over multiple random seeds with standard deviation included.

### v3

@@ 5 Discussion
Efficient, high-performing sparse networks are crucial in resource-constrained environments. However, sparse models cannot easily leverage the benefits of parameter averaging. We addressed this issue proposing SMS, a technique that merges models while preserving sparsity, substantially enhancing IMP and outperforming multiple baselines. By integrating SMS into magnitude-pruning methods during training, we elevated their performance and competitiveness. Despite the focus on pruning, a single type of network compression, we think that our work serves as an important step towards understanding and improving sparsification algorithms.
Acknowledgements
This research was partially supported by the DFG Cluster of Excellence MATH+ (EXC-2046/1, project id 390685689) funded by the Deutsche Forschungsgemeinschaft (DFG) as well as by the German Federal Ministry of Education and Research (fund number 01IS23025B). We would like to thank Berkant Turan and Christophe Roux for providing useful comments.
Appendices
@@ Appendix A Technical details and training settings
A.1 Pretraining
Training settings and metrics.
Table 3 shows the exact pretraining settings for each dataset-architecture pair, reporting the number of epochs used for pretraining, the batch size, weight decay as well as the learning rate used. We stick to SGD as the optimizer, noting that a variety of other optimization methods for training deep neural networks exist (see e.g. Kingma & Ba, 2014; Pokutta et al., 2020). We keep momentum at the default value of 0.9. The last column reports the performance we achieve when performing regular dense training. For image classification tasks, we report the top-1 test accuracy being the fraction of correctly classified test samples. For semantic segmentation, we used pretrained backbones and evaluated the mean Intersection-over-Union (IoU) on the validation dataset, which we use as the test set. For NMT, we report the BLEU score on the test set with sequence length limited to 128. We utilized label smoothing and gradient clipping for MaxViT. If needed, we report the theoretical speedup, a metric indicating the FLOPs ratio for inference between dense and sparse models (Blalock et al., 2020). The speedup, defined as $F_{d}/F_{s}$ where $F_{d}$ and $F_{s}$ are the FLOPs required for dense and pruned models respectively, depends solely on the distribution of pruned weights, not on the values attained by non-zero parameters. FLOPs are computed using a single test batch, with code adapted from the ShrinkBench framework (Blalock et al., 2020).
Table 3: Exact pretraining configurations in our experiments.
@@ A.2 Pruning and Retraining
Identifying which weights to remove is essential for successful magnitude pruning, with multiple methods developed to address this. Zhu & Gupta (2017) presented the Uniform allocation that prunes each layer to the same relative sparsity level. Gale et al. (2019) refined this approach to Uniform+, leaving the first convolutional layer dense and capping pruning in the final fully-connected layer at 80%. Evci et al. (2020) reformulate the Erdős-Rényi kernel (ERK) (Mocanu et al., 2018) to consider layer and kernel dimensions for layerwise sparsity distribution. Lee et al. (2020) suggested Layer-Adaptive Magnitude-based Pruning (LAMP), which minimizes output distortion at pruning time measured by the $L_{2}$-distortion on the worst-case input. Throughout this work, we stick to the Global allocation, in which all trainable parameters are treated as a single vector and a global threshold is computed to remove parameters independent of the layer they belong to. In experiments where we prune convolutional filters instead of weights, we adopt the $L_{2}$-norm criterion from Li et al. (2016), ensuring a uniform sparsity distribution across layers.
@@
We follow the recommendations of Evci et al. (2020) and Dettmers & Zettlemoyer (2019) and refrain from pruning biases and batch-normalization parameters, as their negligible weight contribution is offset by their significant performance impact. Moreover, for GMP experiments, we opt for the global selection criterion due to its superior performance compared to Uniform+ (Zimmer et al., 2023).
Retraining schedules.
The choice of learning rate schedule during the retraining phase has recently attracted interest due to its significant influence on the performance of pruned networks. To avoid the undesired necessity of individually tuning the learning rate schedule in each phase, various retraining schedules have been devised to transpose the original learning rate schedule to the retraining phase. We briefly outline these schedules. Let $T$ represent the total epochs the original network is trained for with the learning rate schedule $(\eta_{t})_{t\leq T}$, and let $T_{{rt}}$ denote the epochs allocated for retraining per prune-retrain cycle. The following retraining schedules have been proposed.
  • Fine Tuning (FT, Han et al., 2015): Retrains the pruned network using the constant learning rate, $\eta_{T}$, from the last epoch of the original training.
  • Learning Rate Rewinding (LRW, Renda et al., 2020): Utilizes the last $T-T_{{rt}}$ learning rates from the original training.
  • Scaled Learning Rate Restarting (SLR, Le & Hua, 2021): Compresses the original learning rate schedule into the retraining timeframe with a short warm-up.
  • Cyclic Learning Rate Restarting (CLR, Le & Hua, 2021): Employs a cosine based schedule with a short warm-up to $\eta_{1}$.
  • Linear Learning Rate Restarting (LLR, Zimmer et al., 2023): A linear decay from $\eta_{1}$ to zero during each retrain cycle.
  • Adaptive Linear Learning Rate Restarting (ALLR, Zimmer et al., 2023): LLR but dynamically adapts the initial learning rate based on the impact of the previous pruning step and the retraining time available, addressing both the length of a prune-retrain cycle and the performance drop induced by pruning.
Hyperparameters for Retraining.
The choice of hyperparameters during retraining significantly impacts the tradeoff between model performance and achieved sparsity. We briefly discuss the hyperparameters applied during the retraining phase, differentiating them from those we leave unchanged compared to the pretraining phase. The exact retraining hyperparameters are specified explicitly in the descriptions of each experiment or in the corresponding subsection in Appendix B.
Specifically, we retain the same batch size and weight decay parameters as used in pretraining. For each IMP-based experiment, we treat the retraining learning rate schedule, the number of retraining epochs, and the number of phases as experiment-specific hyperparameters.
  • Learning Rate Schedule: The retraining schedule has a dramatic impact on the final performance, as outlined in the previous paragraph. Zimmer et al. (2023) demonstrate that LLR and ALLR surpass previously proposed methods across a broad spectrum of architectures, sparsity levels, and retraining durations. Thus, we adhere to these schedules in our experiments, specifying our choice explicitly when important.
  • Number of Retraining Epochs: The number of epochs in retraining influences the extent to which the model can recover the pre-pruned accuracy. Further, in the high sparsity regime, sufficient retraining is required to ensure that models are averageable, as highlighted in 3(b).
  • Number of Phases: The number of prune-retrain phases similarly impacts the performance vs. sparsity tradeoff. In general, high goal sparsity levels require multiple phases.
@@
Throughout our experiments, we also explored variants of LearnedSoup (Wortsman et al., 2022a), which learns the coefficients $\lambda_{i}$ to maximize the validation accuracy. Specifically, we observed improvements when utilizing knowledge distillation techniques (Hinton et al., 2015) with the original pretrained model as the teacher, instead of minimizing the validation loss as suggested by Wortsman et al. (2022b). Nevertheless, these improvements were marginal, so we opted for UniformSoup and GreedySoup for simplicity and to avoid introducing new hyperparameters.
Moreover, we noticed that in later IMP phases, assuming standard IMP rather than starting from an averaged model as in SMS, sparse connectivities tend to diverge, leading to diminished sparsity upon averaging. Specifically, while in the first phase two model splits converge to the same loss basin, subsequent pruning may project them into different subspaces, motivating the combination of individual models into a single one at the end of each phase to ensure that effectively averaging them remains possible. Therefore, we also explored weight alignment, as notably proposed by Ainsworth et al. (2023) and Singh & Jaggi (2020) (see Section 4 for a detailed discussion), hoping to permute models in later phases to a common linear subspace. Although we partially mitigated the sparsity reduction, we were unable to fully recover the original sparsity. This approach may only work if different IMP runs share the same distribution of sparsity among layers (i.e., a specific layer in model one must have the same sparsity as the same layer in model two), which is generally not the case.
@@ A.3 Pruning during Training
Appendix B Extended results
This section contains additional tables and plots. The subsections follow the same structure as the main experimental section. To maintain transparency, we explicitly mention the retraining schedule and duration in the captions of tables and figures, or in the beginning of the subsection if suitable.
@@ B.1 Evaluating Sparse Model Soups
Table 4: WideResNet-20 on CIFAR-100 (unstructured pruning): Test accuracy comparison of SMS to several baselines for target sparsities 90% (top) and 98% (bottom) given three prune-retrain cycles. We report results using UniformSoup as well as GreedySoup for merging, employing ALLR as the retraining schedule for 10 epochs of retraining per phase. Results are averaged over multiple seeds with standard deviation included. The best value is highlighted in bold.
@@
Table 5: ResNet-18 on CIFAR-10 (unstructured pruning): Test accuracy comparison of SMS to several baselines for target sparsity 98% given three prune-retrain cycles. We report results using UniformSoup as well as GreedySoup for merging, employing ALLR as the retraining schedule for 20 epochs of retraining per phase. Results are averaged over multiple seeds with standard deviation included. The best value is highlighted in bold.
@@
Table 6: ResNet-18 on CIFAR-10 (unstructured pruning): Test accuracy comparison of SMS to several baselines for target sparsities 80%, 90%, 95% in the One Shot setting, using ALLR for a retrain length of 20 epochs per phase. We report results using UniformSoup as well as GreedySoup for merging. Results are averaged over multiple seeds with standard deviation included. The best value is highlighted in bold.
@@
Table 7: MaxViT on ImageNet (unstructured pruning): Test accuracy comparison of SMS to several baselines for target sparsities 75%, 80%, 85% in the One Shot setting, using ALLR for a retrain length of 10 epochs per phase. We report results using UniformSoup as well as GreedySoup for merging. Results are averaged over multiple seeds with standard deviation included. The best value is highlighted in bold.
@@
Table 8: PSPNet on Cityscapes (unstructured pruning): Test accuracy comparison of SMS to several baselines for target sparsity 90% given two prune-retrain cycles of 50 retraining epochs each. We report results using UniformSoup as well as GreedySoup merging. Results are averaged over multiple seeds with standard deviation included. The best value is highlighted in bold.
@@
SMS (uniform) 58.52 ±0.15 58.47 ±0.10 58.73 ±0.20 58.40 ±0.30
best candidate 58.20 ±0.37 58.25 ±0.48 58.62 ±0.60 57.92 ±0.36
mean candidate 57.96 ±0.29 57.80 ±0.33 58.38 ±0.41 57.62 ±0.48
SMS (greedy) 58.14 ±0.14 58.63 ±0.36 58.46 ±0.24 58.79 ±0.09
best candidate 58.26 ±0.06 58.59 ±0.14 58.13 ±0.27 58.73 ±0.06
mean candidate 57.82 ±0.13 58.17 ±0.14 57.90 ±0.09 58.15 ±0.04
$\text{IMP}_{m\!\times}$ 57.39 ±0.06 58.35 ±0.08 58.27 ±0.42 58.39 ±0.17
IMP-RePrune — N/A — 58.10 ±0.24 58.64 ±0.34
IMP — 57.92 ±0.04 — 58.89 ±0.23
Table 9: PSPNet on Cityscapes (unstructured pruning): Test accuracy comparison of SMS to several baselines for target sparsities 60%, 70%, 80% and 90% in the One Shot setting, using LLR as the retraining schedule for 50 epochs of retraining. We report results using UniformSoup as well as GreedySoup merging. Results are averaged over multiple seeds with standard deviation included. The best value is highlighted in bold.
@@
SMS (uniform) 58.11 ±0.22 57.59 ±0.38 58.40 ±0.03 58.19 ±0.28
best candidate 57.74 ±0.11 57.36 ±0.43 57.97 ±0.06 57.87 ±0.49
mean candidate 57.37 ±0.47 57.04 ±0.48 57.88 ±0.03 57.70 ±0.44
SMS (greedy) 58.41 ±0.13 58.13 ±0.31 57.95 ±0.40 57.30 ±0.21
best candidate 58.16 ±0.49 57.78 ±0.13 57.78 ±0.26 57.32 ±0.28
mean candidate 58.05 ±0.50 57.55 ±0.07 57.49 ±0.38 57.18 ±0.25
$\text{IMP}_{m\!\times}$ 58.02 ±0.09 58.09 ±1.04 58.26 ±0.13 58.47 ±0.22
IMP 57.44 ±0.71 58.05 ±0.28 57.43 ±0.24 56.99 ±0.69
Table 10: T5 on WMT16 (unstructured pruning): BLEU score comparison of SMS to several baselines for target sparsities 50%, 60%, 70% in the One Shot setting. We report results using UniformSoup as well as GreedySoup merging, employing ALLR as the retraining schedule for 2 epochs of retraining per phase. Results are averaged over multiple seeds with standard deviation included. The best value is highlighted in bold.
@@
SMS (greedy) 25.51 ±0.28 24.92 ±0.47 24.14 ±0.02
best candidate 25.39 ±0.03 24.96 ±0.26 24.12 ±0.01
mean candidate 25.16 ±0.08 24.79 ±0.19 24.03 ±0.01
$\text{IMP}_{m\!\times}$ 25.36 ±0.12 25.09 ±0.05 24.00 ±0.04
@@
Table 11: WideResNet-20 on CIFAR-100 (structured pruning): Test accuracy comparison of SMS to several baselines for target sparsities 60% (top) and 80% (bottom) given three prune-retrain cycles. We report results using UniformSoup as well as GreedySoup for merging, employing ALLR as the retraining schedule for 10 epochs of retraining per phase. Results are averaged over multiple seeds with standard deviation included. The best value is highlighted in bold.
@@
Table 12: ResNet-18 on CIFAR-10 (structured pruning): Test accuracy comparison of SMS to several baselines for target sparsities 40%, 50%, 60% in the One Shot setting, using ALLR as the retrain schedule for a retrain length of 20 epochs. We report results using UniformSoup as well as GreedySoup for merging. Results are averaged over multiple seeds with standard deviation included. The best value is highlighted in bold.
@@ B.2 Examining Sparse Model Merging
B.2.1 Exploring parameters beyond random seeds.
@@
To generate the averaging candidates, we employed the following hyperparameter configurations. For ImageNet, as showcased in Figure 3, our base configuration utilized ALLR for 5 retraining epochs with weight decay as stated in Table 3. While varying the random seed, we maintained the base configuration and selected eight distinct random seeds. In adjusting the weight decay, we adhered to the base configuration and experimented with weight decay strengths of 4e-5, 6e-5, 8e-5, 1e-4, 1.2e-4, 1.4e-4, 1.6e-4, 1.8e-4. For retraining length variation, we examined all integral values between 2 and 9 epochs. In terms of retraining schedule modification, we adopted a linearly decaying learning rate schedule, tuning the initial value among 2e-2, 4e-2, 6e-2, 8e-2, 1e-1, 1.2e-1, 1.4e-1 and 1.6e-1. For CIFAR-100, as in Figure 5, we used ALLR for 10 epochs with weight decay as in Table 3. To adjust the weight decay, we experimented with weight decay strengths of 1e-4, 2e-4, 3e-4, 4e-4, 5e-4, 6e-4, 7e-4, 8e-4. For retraining length variation, we examined all integral values between 6 and 13 epochs. In terms of retraining schedule modification, we adopted a linearly decaying learning rate schedule, tuning the initial value among 6e-2, 7e-2, 8e-2, 9e-2, 1e-1, 1.1e-1, 1.2e-1 and 1.3e-1.
@@
B.2.2 Instability to randomness and recovering it.
Figure 6 replicates the plots from Figure 4, albeit for CIFAR-10, adhering to the identical retraining hyperparameter configuration delineated in the main text.
@@
Figure 6: ResNet-18 on CIFAR-10: (a) Accuracy difference between the soup and best performing model after One Shot pruning and retraining. The lines for UniformSoup and GreedySoup show the envelope considering all retraining schedules and durations. (b) Accuracy difference between the soup ($m=3$) and $\text{IMP}_{3\!\times}$ retrained three times as long as indicated on the x-axis, using One Shot pruning to 90% and 98% sparsity. Results are averaged over multiple random seeds with min-max bands indicated.
B.2.3 OOD-Robustness of Sparse Model Soups.
Figure 7 depicts OOD-robustness effects for One Shot IMP on WRN-20 trained on CIFAR-100 at 90% sparsity, while Figure 8 does the same for ResNet-50 on ImageNet at 70% sparsity. For each parameter (random seed, weight decay strength, retraining duration, and initial learning rate in a linear decay schedule), we formulated eight variations centered on default values (see subsubsection B.2.1 for exact hyperparameters). Contrary to previous scatter plots, these evaluate models on the robustness benchmarks CIFAR-100-C and Imagenet-C. The plots contrast the OOD accuracy in the uniform soup setting (averaged across model pairs) with the peak OOD accuracy of each pair. The OOD accuracy is computed on the entire corrupted dataset, i.e., among all corruption types and degrees of severity (ranging from 1 to 5).
Further, Table 13 and Table 14 display the OOD-robustness evaluated on CIFAR-100-C and ImageNet-C, respectively, when aiming for higher sparsities and using multiple cycles. SMS consistently improves over the baselines and improves the out-of-distribution accuracy significantly. For the pretrained models we obtain a base ood accuracy of $49.03\%(\pm 0.33)$ for CIFAR-100-C and $40.35\%(\pm 0.30)$ for ImageNet-C.
Figure 7: WideResNet-20 evaluated on CIFAR-100-C: OOD Accuracy of average of two models vs. the maximal individual OOD accuracy. All models are pruned to 90% sparsity (One Shot) and retrained, varying the indicated hyperparameters.
Figure 8: ResNet-50 evaluated on ImageNet-C: OOD Accuracy of average of two models vs. the maximal individual OOD accuracy. All models are pruned to 70% sparsity (One Shot) and retrained, varying the indicated hyperparameters.
Table 13: WideResNet-20 trained on CIFAR-100 and evaluated on CIFAR-100-C (unstructured pruning): OOD accuracy comparison of SMS to several baselines for target sparsities 90% (top) and 98% (bottom) given three prune-retrain cycles. We only report results using UniformSoup for merging, employing ALLR as the retraining schedule for 10 epochs of retraining per phase. Results are averaged over multiple seeds with standard deviation included. The best value is highlighted in bold.
CIFAR-100 (90%)
Accuracy of $m=3$ $m=5$ $m=10$ $m=3$ $m=5$ $m=10$ $m=3$ $m=5$ $m=10$
SMS 49.60 ±0.14 49.48 ±0.20 49.57 ±0.24 49.90 ±0.11 50.09 ±0.08 50.33 ±0.14 48.65 ±0.31 48.86 ±0.06 49.38 ±0.14
best candidate 49.31 ±0.14 49.18 ±0.22 49.17 ±0.23 48.51 ±0.09 48.64 ±0.08 48.82 ±0.17 47.42 ±0.29 47.87 ±0.15 47.87 ±0.09
mean candidate 49.10 ±0.18 48.97 ±0.15 48.97 ±0.18 48.40 ±0.12 48.33 ±0.11 48.37 ±0.03 47.29 ±0.28 47.51 ±0.20 47.61 ±0.07
$\text{IMP}_{m\!\times}$ 49.25 ±0.10 49.01 ±0.52 49.17 ±0.38 48.06 ±0.07 48.24 ±0.52 48.17 ±0.65 47.10 ±0.74 46.59 ±0.05 47.12 ±0.52
IMP-RePrune — N/A — — N/A — 48.72 ±0.25 49.03 ±0.15 49.35 ±0.11
IMP — 49.41 ±0.22 — — 48.25 ±0.08 — — 46.88 ±0.27 —
CIFAR-100 (98%)
Accuracy of $m=3$ $m=5$ $m=10$ $m=3$ $m=5$ $m=10$ $m=3$ $m=5$ $m=10$
SMS 50.05 ±0.09 50.01 ±0.07 50.28 ±0.09 48.30 ±0.15 48.52 ±0.07 48.80 ±0.12 44.18 ±0.43 44.81 ±0.26 44.80 ±0.75
best candidate 49.36 ±0.17 49.01 ±0.16 49.18 ±0.02 47.11 ±0.12 47.07 ±0.20 47.27 ±0.25 43.37 ±0.41 43.49 ±0.23 43.72 ±0.64
mean candidate 48.95 ±0.07 48.68 ±0.09 48.67 ±0.09 46.87 ±0.14 46.80 ±0.08 46.87 ±0.23 42.94 ±0.52 43.31 ±0.34 43.17 ±0.54
$\text{IMP}_{m\!\times}$ 48.70 ±0.19 48.81 ±0.10 48.66 ±0.17 46.20 ±0.27 45.90 ±0.10 46.17 ±0.07 41.97 ±0.17 41.62 ±1.55 42.76 ±0.30
IMP-RePrune — N/A — — N/A — 39.57 ±0.82 37.17 ±1.04 35.28 ±1.26
IMP — 48.60 ±0.14 — — 45.89 ±0.14 — — 42.43 ±0.58 —
Table 14: ResNet-50 trained on ImageNet and evaluated on ImageNet-C (unstructured pruning): OOD accuracy comparison of SMS to several baselines for target sparsity 90% given three prune-retrain cycles. We only report results using UniformSoup for merging, employing ALLR as the retraining schedule for 10 epochs of retraining per phase. Results are averaged over multiple seeds with standard deviation included. The best value is highlighted in bold.
ImageNet (90%)
Accuracy of $m=3$ $m=5$ $m=10$ $m=3$ $m=5$ $m=10$ $m=3$ $m=5$ $m=10$
SMS 42.19 ±0.17 42.65 ±0.17 42.94 ±0.02 41.17 ±0.15 41.62 ±0.10 42.11 ±0.06 38.70 ±0.02 39.23 ±0.15 39.70 ±0.02
best candidate 39.98 ±0.12 40.00 ±0.25 40.11 ±0.05 39.31 ±0.11 39.43 ±0.03 39.66 ±0.05 37.30 ±0.01 37.57 ±0.13 37.84 ±0.11
mean candidate 39.87 ±0.10 39.90 ±0.22 39.91 ±0.09 39.19 ±0.10 39.28 ±0.05 39.47 ±0.01 37.21 ±0.01 37.39 ±0.14 37.62 ±0.03
$\text{IMP}_{m\!\times}$ 40.05 ±0.17 40.36 ±0.29 40.44 ±0.11 39.31 ±0.02 39.46 ±0.05 39.45 ±0.28 37.13 ±0.22 37.47 ±0.26 37.36 ±0.34
IMP-RePrune — N/A — — N/A — 37.10 ±0.15 36.81 ±0.18 36.48 ±0.66
IMP — 39.84 ±0.05 — — 38.74 ±0.02 — — 36.64 ±0.00 —
B.2.4 Reducing compression-induced unfairness.
@@
Classification model performance, usually measured by the top-1 accuracy, can mask the disproportionate effect of compression on individual class performance (Hooker et al., 2019; 2020; Paganini, 2020). Pruning often sacrifices difficult samples, benefiting well-performing classes and worsening the performance of others (Tran et al., 2022). Recent research by Ko et al. (2023) highlights the benefits of prediction ensembling in enhancing fairness metrics, including minority group performance.
@@
Table 15: ResNet-18 on Celeb-A: Comparison of the pretrained (i.e. dense) base model against SMS and IMP for different sparsity levels, employing ALLR as the retraining schedule for 20 epochs of retraining. We indicate the top-1 test accuracy as well as the recall on the four different sub-groups. All results are averaged over multiple random seeds with standard deviation included.